# Why Is Spatial Reasoning Hard for VLMs? An Attention Mechanism Perspective on Focus Areas

### v2

@@ Abstract
Shiqi Chen1Tongyao Zhu5Ruochen Zhou1Jinghan Zhang4Siyang Gao1Juan Carlos Niebles2,7
Mor Geva6Junxian He4Jiajun Wu2Manling Li2,3
1City University of Hong Kong  2Stanford University  3Northwestern University
4Hong Kong University of Science and Technology  5National University of Singapore  6Tel Aviv University
7Salesforce Research
@@ 2 Preliminary on VLMs
where attention logits ${\bm{A}}^{(l,h)}$ are computed via ${\bm{Q}}={\bm{X}}{\bm{W}}_{q_{h}},{\bm{K}}={\bm{X}}{\bm{W}}_{k_{h}},{\bm{V}}={\bm{X}}{\bm{W}}_{v_{h}}$, and ${\bm{W}}_{q_{h}},{\bm{W}}_{k_{h}},{\bm{W}}_{v_{h}}\in\mathbb{R}^{d\times d_{h}}$ are learnable projection matrix of the head $N_{h}$. A causal mask ${\bm{M}}_{ij}=0$ if $i\geq j$, and $\infty$- ∞ otherwise, prevents tokens from attending to future tokens.
@@ 3 Text-Vision Attention Interactions
[…]
@@ 5.2 AdaptVis
Table 1: Results on Controlled A, Controlled B, COCO, and VG datasets. (Metrics in ×$\times 10^{-2}$). Best-performing method per model and dataset are highlighted in bold. P Acc and S Acc represents Pair Acc and Set Acc.
@@
Table 2: Results on the VSR dataset (Exact Match and F1) (Metrics in ×$\times 10^{-2}$). Best-performing method per model and dataset are highlighted in bold.
@@ 7 Conclusion and Future Work
Appendix
[…]
Appendix A Limitations
Appendix B Broad Impact
@@ E.1 Other metrics to distinguish the distributions
In Equation 4,$\mathcal{A}_{n,j}^{(l,h)}$ denotes the attention scores assigned by the $h$-th head in the $l$-th layer to the $j$-th token in sequence $n$. The summation runs over $j=1$ to $t$, where $t$ is the total number of tokens considered for this attention distribution. $\tilde{P}\left(\mathcal{A}_{n,j}^{(l,h)}\right)$ is the normalized probability distribution of these attention scores. This entropy measures the uncertainty or spread of the attention distribution across tokens.
@@ E.3 Hyperparameter choice and robustness
Table 4: OOD test results on WhatsUp (Metrics in ×$\times 10^{-2}$) for LLaVA1.5. Arrows show growth over baseline. P-Acc and S-Acc are Pair and Set Accuracy.
@@ E.6 What changes after our intervention?
How the confidence changes by using different $\alpha$s?
@@ E.8 Reverse curse?
Best $\alpha$ $0.5$ $0.5$ $2$ $2$
Figure 27: Results on WhatsUp (Metrics in ×$\times 10^{-2}$) when changing prompts for other option numbers. Arrows show improvement over greedy decoding.

### v3

@@ Abstract
Machine Learning, ICML
@@ 2 Preliminary on VLMs
where attention logits ${\bm{A}}^{(l,h)}$ are computed via ${\bm{Q}}={\bm{X}}{\bm{W}}_{q_{h}},{\bm{K}}={\bm{X}}{\bm{W}}_{k_{h}},{\bm{V}}={\bm{X}}{\bm{W}}_{v_{h}}$, and ${\bm{W}}_{q_{h}},{\bm{W}}_{k_{h}},{\bm{W}}_{v_{h}}\in\mathbb{R}^{d\times d_{h}}$ are learnable projection matrix of the head $N_{h}$. A causal mask ${\bm{M}}_{ij}=0$ if $i\geq j$, and -$\infty$ otherwise, prevents tokens from attending to future tokens.
@@ 3 Text-Vision Attention Interactions
For evaluation, we use LLaVA-1.5 (Liu et al., 2024a) throughout the analysis part (Section 3 and 4). For metric, we apply accuracy of exact match as the primary metric. To maintain consistency in the label space across datasets, we use a four-option setting $\langle\textit{left},\textit{right},\textit{on},\textit{under}\rangle$ for the Controlled_Image and COCO subsets, and a six-option setting $\langle\textit{left},\textit{right},\textit{on},\textit{under},\textit{behind},\textit{front}\rangle$ for the VG as it contains additional spatial annotations.
@@ 5.2 AdaptVis
Table 1: Results on Controlled A, Controlled B, COCO, and VG datasets. (Metrics in $\times 10^{-2}$). Best-performing method per model and dataset are highlighted in bold. P Acc and S Acc represents Pair Acc and Set Acc.
@@
Table 2: Results on the VSR dataset (Exact Match and F1) (Metrics in $\times 10^{-2}$). Best-performing method per model and dataset are highlighted in bold.
@@ 7 Conclusion and Future Work
Impact Statement
@@
Acknowledgments
We thank Min Lin, Miao Xiong, Guangtao Zeng, Teng Xiao, Wei Liu, Chi Han, Chang Ma, Zhengxuan Wu, and Junlong Li for their valuable discussions and insights.
Appendix
Appendix A Limitations
@@
Appendix B Additional Results on Qwen2-VL
We further expand our intervention experiments and attention analysis on Qwen2-VL, the results on spatial reasoning benchmarks and on general benchmarks are shown in Table 3 and Table 4 respectively. We intervene in the image attention distribution using our temperature-scaling method, showing consistent improvements below, particularly in challenging cases. For example, on VG_two_object, where the baseline performance is the lowest among all benchmarks, our method yields a significant improvement of 10+ absolute points. The gains observed on Qwen2-VL further demonstrate the generalizability of our approach. Experiments on more benchmarks shown in Table 4, including POPE (Li et al., 2023), GQA (Hudson & Manning, 2019), and TextVQA (Singh et al., 2019), which proves that attention intervention can maintain the performance on more general tasks without hurting the performance. Compared with spatial reasoning tasks, general QA tasks achieve relatively smaller improvement. A possible reason is that such tasks are less sensitive to the geometric structured distribution of image attention. For example, given a question like “Is there a dog in this picture?”, the model only needs to detect the presence of an object, and is therefore less likely to suffer from misallocated attention across spatial regions.
Benchmark Qwen2-VL +Attention Intervention
VSR 78.96 81.60 $\uparrow$2.64
Coco_one_obj 76.64 78.03 $\uparrow$1.39
Coco_two_obj 75.28 76.52 $\uparrow$1.24
VG_one_obj 74.89 75.11 $\uparrow$0.22
VG_two_obj 56.22 66.95 $\uparrow$10.73
Controlled_A 98.18 98.18 $\uparrow$0.00
Controlled_B 91.73 92.97 $\uparrow$1.24
Table 3: Performance of Qwen2-VL before and after the attention intervention on spatial reasoning benchmarks.
Benchmark Qwen2-VL +Attention Intervention
POPE-Overall 86.32 87.09 $\uparrow$0.77
POPE-P 86.47 87.29 $\uparrow$0.82
POPE-A 85.07 85.80 $\uparrow$0.73
POPE-R 87.46 88.22 $\uparrow$0.76
GQA 62.09 62.17 $\uparrow$0.08
TextVQA 79.18 79.26 $\uparrow$0.08
Table 4: Performance of Qwen2-VL before and after the attention intervention on additional general benchmarks.
We also conduct an attention analysis to verify whether our observation of the sparsity of image attention compared to textual tokens still holds. The results are shown in Table 6. Here we calculate the Qwen2-VL’s attention scores below (average attention logits for single text/image tokens respectively), which matches our previous claim that image receives much less attention than the text tokens is generally valid for VLMs. And we also conduct the uncertainty experiments to show our uncertainty conclusion is also valid. The results are shown in Table 5. It shows that spatial relationships with lower confidence, such as “Left” and “Under” tend to exhibit lower accuracy when compared to those with higher confidence, like “On” and “Right”. After the intervention with AdaptVis, certain spatial relationships like “Left” show improved performance, accompanied by an increase in confidence. This demonstrates a pattern consistent with our observations for LLaVA, as depicted in Figure 8 in Section §4.
Benchmark Text attention Image attention
Controlled_A 1.57e-02 7.59e-05
Controlled_B 1.58e-02 7.69e-05
Coco_one_obj 1.77e-02 4.48e-04
Coco_two_obj 1.65e-02 3.50e-04
VG_one_obj 1.54e-02 4.75e-04
VG_two_obj 1.42e-02 4.19e-04
Table 5: Average text-token and image-token attention scores across benchmarks.
Benchmark Qwen2‑VL Unc. +ScalingVis Unc. Qwen2‑VL Acc +ScalingVis Acc
Left 0.4408 0.4474 70.11 74.71
On 0.5758 0.5668 100 100
Right 0.5982 0.5911 100 100
Under 0.5554 0.5418 98.82 98.82
Table 6: Uncertainty (Unc.) and accuracy (Acc) on spatial‑relation categories before and after applying ScalingVis.
@@ E.1 Other metrics to distinguish the distributions
[…]
@@ E.3 Hyperparameter choice and robustness
Table 8: OOD test results on WhatsUp (Metrics in $\times 10^{-2}$) for LLaVA1.5. Arrows show growth over baseline. P-Acc and S-Acc are Pair and Set Accuracy.
@@ E.6 What changes after our intervention?
How the confidence changes by using different $\alpha$?
@@ E.8 Reverse curse?
Best $\alpha$ $0.5$ $0.5$ 22 22
Figure 27: Results on WhatsUp (Metrics in $\times 10^{-2}$) when changing prompts for other option numbers. Arrows show improvement over greedy decoding.